### v1

# Membership Inference Attacks on LLM-based Recommender Systems

Abstract
Recommender systems (RecSys) have become an essential component of many web applications. The core of the system is a recommendation model trained on highly sensitive user-item interaction data. While privacy-enhancing techniques are actively studied in the research community, the real-world model development still depends on minimal privacy protection, e.g., via controlled access. Users of such systems should have the right to choose not to share highly sensitive interactions. However, there is no method allowing the user to know which interactions are more sensitive than others. Thus, quantifying the privacy risk of RecSys training data is a critical step to enabling privacy-aware RecSys model development and deployment. We propose a membership-inference attack (MIA)- based privacy scoring method, RecPS, to measure privacy risks at both the interaction and user levels. The RecPS interaction-level score definition is motivated and derived from differential privacy, which is then extended to the user-level scoring method. A critical component is the interaction-level MIA method RecLiRA, which gives high-quality membership estimation. We have conducted extensive experiments on well-known benchmark datasets and RecSys models to show the unique features and benefits of RecPS scoring in risk assessment and RecSys model unlearning.
1 Introduction
Recommendation systems (RecSys) have seen significant advances over the past decade and have been widely used in various scenarios such as job matching [11], e-commerce [10], and entertainment [18]. However, one critical challenge is still difficult to tackle: recommendation models are naturally task-specific, as they are typically trained with task-specific user-item interactions [30, 44]. Therefore, it is almost impossible to move a recommendation system developed for one domain to another without significant performance degradation. Collecting new training data to build a new recommendation system has been a well-accepted practice. However, it is expensive and time-consuming. Practitioners and researchers have been looking for more efficient approaches to addressing this domain locked-in issue.
With the emerging applications of large language models (LLMs), researchers wonder whether LLMs can be used to gain the highly desired low-cost cross-domain generalization power of RecSys. Earlier efforts focused on fine-tuning general-purpose LLMs for a specific recommendation domain, including P5 [19], M6-Rec [14], and TALLRec [3], which often involved expensive adjustments. Most recent studies turned to In-Context Learning (ICL) methods that use cheap zero-shot or few-shots prompts to simplify the customization of LLM for RecSys [30, 25]. These methods have shown comparable or even better recommendation quality [25, 44] than more expensive and sophisticated fine-tuning methods.
While ICL offers substantial advantages through the power of prompts, its integration into language models raises a critical issue: privacy leakage via prompts [39]. It is well known that the success of recommendation systems is based on user-interaction data, which often contains private information about user preferences, activities, and social contexts [9, 42, 46, 38]. In few-shot ICL RecSys, such data is used by RecSys to design personalized system prompts. Figure 1 shows a typical example, where the user’s sensitive historical interactions marked in red.
Figure 1: Prompt Example for LLM RecSys
One of the most fundamental privacy attacks is the membership inference attack (MIA) [5, 27, 22] that identifies whether a record is used in the model training dataset. While most MIAs focused on classification modeling, researchers have recently identified the unique features of traditional RecSys models for MIAs, such as matrix factorization RecSys [24], federated RecSys [28], and knowledge-based RecSys [37], and designed several RecSys-specific MIA methods [42, 41, 46, 23]. However, LLM-based RecSys have several unique features that the MIA methods designed for traditional RecSys models cannot be directly applied.
  • The traditional RecSys MIAs utilize the system output, i.e., the recommended items, to infer whether the core RecSys machine learning model has used the target user’s information in training. The inference utilizes the similarity between the recommended items and the known victim user’s interacted items, where item embedding, i.e., item’s vector representation, is used in the similarity evaluation. The item embedding is generated from a large number of existing user-item interactions, via, e.g., matrix factorization [24] methods, which works effectively for non-LLM RecSys. However, this is a very strong attack assumption. In practice, the interaction matrix may not be accessible. We do not know whether and how the similarity-based method still works for LLM-based RecSys outputs, e.g., via general text semantic embedding.
  • Existing RecSys MIAs [43, 41, 46] assume that the training data distribution is known by the adversary, which is used to train shadow models offline to mimic the behavior of the target model, which are then used to build the attack model. In LLM-based RecSys, the training data now appear in the system prompts. The concept of shadow models needs to be re-examined and it is unclear whether the assumption of known training data distribution is necessary and how it can be used in attacks.
  • LLMs have some distinct features that other machine learning models do not have, such as hallucination [25] and memorization [6]. These features might enable new attacks distinct from those on traditional RecSys models. However, no study has been done on attacks utilizing such features for LLM RecSys.
To our knowledge, no MIA method has been reported on LLM-based RecSys. Understanding such new MIA methods will allow LLM RecSys designers to be aware of potential privacy threats and incorporate privacy protection measures into system design.
Scope of Our Research. We design, evaluate and analyze four membership inference attacks on LLM-powered RecSys that use in-context learning, i.e., prompts, to customize the recommendation. These attacks target the private user-item interactions [^1] In the following, for simplicity, we use “interactions” to represent tuples (user, item, interaction score). embedded in system prompts by the LLM RecSys provider. We assume that the attacker has known the target user’s historical interactions but not known whether any of these interactions were used by the LLM RecSys to compose the system prompts.
  • Direct Inquiry attack utilizes the basic generative features of LLMs, where the attacker directly asks the LLM whether it has seen the victim user.
  • Hallucination attack investigates the likelihood that LLM hallucination may occur more often when the system has not seen the query. This attack explores the novel application of the unique LLM hallucination phenomenon.
  • Similarity attack identifies target users as members if items recommended by LLM have high similarity to the user’s historical interactions known by the adversary. It resembles the similarity attack in the traditional RecSys MIA. However, since the RecSys-specific item embedding is not available for LLMs, we will depend on general text embedding to evaluate similarity.
  • Poisoning attack uses the modified historical interactions as user-provided additional prompts and observes the change in recommended items to infer whether the target user is a member. The intuition is that if the user’s interactions were used by the system, it may memorize them and the poisoned interactions less likely change the system’s recommendations. This attack explores another unique feature: LLM memorization, in MIAs.
We have conducted extensive experiments on three popular large language models that have been used to develop LLM RecSys and two well-known benchmark datasets: MovieLens-1M and Amazon Digital Music. Empirical results confirm that several attacks work surprisingly well, raising significant alarms for LLM RecSys practitioners. The Inquiry attack achieves over 99% attack advantage, which defined as 2*(MIA accuracy-0.5)*100%, in inferring membership status against Llama3-8b on both datasets; the Poisoning Attack achieves over 70% attack advantage in inferring membership status against Llama2-7b on the MovieLens-1M and over 80% advantage on Amazon Digital Music.
We further analyze factors influencing successful attacks, including the number of shots used by system prompts and the positions of the attacked shots in the prompt. The results shows these factors have varying effect on different LLMs and datasets. These findings may offer insights for designing system prompts that are more resilient against privacy attacks.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to propose and study membership inference attacks against ICL-LLM-powered RecSys.
  • We have designed Direct Inquiry, Hallucination, Similarity, and Poisoning attacks, aiming to effectively detect whether users are used in RecSys system prompts.
  • We have conducted extensive experiments to show the performance of these attacks, among which Direct Inquiry and Poisoning attacks work best for some LLMs. We have also studied the factors affecting these attacks.
The remaining sections include a discussion on the related work focusing on MIAs on RecSys and LLMs (Section 2), the preliminaries about the proposed methods describing how ICL-LLM-powered RecSys works (Section 3), the new threat model (Section 4), the four MIA methods (Section 5), the experimental evaluation exploring the attack performance and factors affecting these attacks (Section 6), and a conclusion about this study (Section 7).
2 Related Work
In this section, we introduce current status of MIA studies on LLM and on RecSys, respectively.
2.1 MIA on LLMs
Membership inference attack (MIA) is one of the most fundamental forms of privacy attacks [5, 27], where an adversary seeks to determine whether a particular sample was part of a model’s training dataset [5, 27]. While widely studied in traditional machine learning [27, 32, 22, 42], MIA has become increasingly concerning in the context of large language models (LLMs), as revealing whether specific data appears in prompts can lead to breaches of sensitive or private information.
Theoretical foundations of MIA largely rely on the observation that models behave more confidently on samples seen during training [5, 27]. A common attack strategy is to train a classifier attack model using the target model’s output posteriors’ probabilities, where higher-confidence outputs are deemed more likely to be members. To improve attack performance, researchers have also incorporated additional cues such as intermediate representations [33], loss trajectories [31], or trained shadow models on crafted datasets [5]. In all of these instances, it seems that having access to the model posterior is a necessary requirement for launching the attack. Most existing membership inference attacks against LLMs necessitate, at a minimum, access to the probability associated with predictions. This requirement is crucial for calculating corresponding loss [16, 40] or perplexity [8, 7], which can then be used to extract membership signals.
Recent work has explored posterior-free MIA techniques [13, 29] that infer membership by estimating a sample’s distance to the decision boundary. However, such approaches have their own challenges due to their black-box nature and the discrete input space. Rui et al. [39] proposed text-only MIAs against LLMs on the classification task which cannot apply to recommendation task due to the entirely different problem settings.
2.2 MIA on RecSys
The earlier RecSys MIA studies are focused on the user level. Zhang et al. [42] propose the Item-Diff method for inferring membership in a target RecSys by analyzing the similarity between a user’s historical interactions and recommended items. The core idea is that, for users in the training set, their historical interactions are likely to be more closely aligned with the items recommended by the system. Wang et al. [38] propose the DL-MIA framework to improve Item-Diff with a VAE-based encoder and weight estimator to address issues with Item-Diff. More recently, Wei et al. [41] proposed a white-box interaction-level membership inference on federated RecSys. Zhong et al. [46] proposed another interaction-level membership inference on Knowledge Graph-based RecSys, utilizing the similarity matrix between the interacted items and the recommended items. To our knowledge, no MIA study is reported on LLM RecSys.
3 Preliminaries
In this section, we introduce In-context learning RecSys in Section 3.1
3.1 In-context learning(ICL) RecSys
In-Context Learning (ICL) is a distinctive feature of large language models (LLMs) [4], enabling them to perform specific tasks by observing a few examples—without updating model parameters. ICL LLMs learn through analogy [15] by appending input-output demonstrations, known as prompts, during inference. Introduced alongside GPT-3 [4],
Figure 2: System Architecture for ICL-RecSys
ICL has proven effective in adapting LLMs to various downstream tasks, particularly in recommendation systems (RecSys). According to Gao et al. [17], its success stems from the design of prompts and in-context demonstrations. In other words, the key innovation of ICL is to elicit the in-context ability of LLMs for learning (new or unseen) downstream tasks from context during the inference stage. In particular, two settings proposed in ICL are prevalently leveraged for prompting LLMs for RecSys: few-shot, where the model is guided by a few input-output examples, and zero-shot, where only natural language task descriptions are provided. Both rely on the model’s ability to generalize from context, but few-shot ICL typically achieves better performance due to the inclusion of in-context demonstrations. Several studies compare these settings under the same recommendation tasks [30, 44, 45] and find that few-shot learning outperforms zero-shot approaches.
To adapt large language models (LLMs) to recommendation tasks via in-context learning (ICL), a straightforward strategy is to prompt LLMs to act as recommenders. For instance, Liu et al. [30] utilize ChatGPT with task-specific prompts for top-k recommendation and explanation generation. Their method involves constructing tailored input-output examples for each task, demonstrating that few-shot settings consistently yield better performance than zero-shot ones. Similarly, other studies have proposed distinct strategies for designing effective in-context demonstrations to enhance recommendation performance. For example, [45] introduces role-based textual descriptions, such as “You are a book rating expert,” to augment in-context prompts. This role injection technique helps mitigate cases where LLMs refuse to perform recommendation tasks (e.g., responding with “As a language model, I don’t have the ability to recommend …”). In summary, ICL serves as a critical bridge connecting LLMs with downstream tasks such as recommendation systems, enabling LLMs to emulate the behavior of traditional recommenders and effectively interact with users.
Figure 2 shows how an LLM-based RecSys may look like. The core components include the LLM, the database containing many users’ historical interactions, and a prompt composer that can adapt to the recommendation task and each specific user’s preferences.
In this work, we propose a text-only membership inference attack on ICL-based RecSys, which operates under the most constrained setting: the adversary only has access to the user interacted item and generated recommended item, not the probabilities, which aligns with the common black-box setting in the recommendation task [42].
Figure 3: Prompt Example
4 THREAT MODEL
4.1 Adversary’s Objective
The primary objective of the adversary is to determine whether a specific target user $u$ was included in the construction of a prompt used to customize a language model $\mathcal{M}$. The prompt, denoted as prompt, comprises a set of $k$ examples, formatted as:
$\text{prompt}=\{\text{Instruction task},(u_{1},i_{1},s_{1}),\dots,(u_{k},i_{k},s_{k})\},$
where $u_{j}$ and $i_{j}$ are from the user set $U$, and item set $I$, correspondingly. $s_{j}$ is a score defined by a specific scoring method, i.e., binary 0/1 for “clicked” or “not clicked” or a multi-scale grade, e.g., from like to dislike. For simplicity, we have used the binary samples – $(u_{k},i_{k})$ implies the user $u_{k}$ interacted $i_{k}$. The detailed prompt sample is shown in Figure 3. The adversary’s goal is to determine whether the target user $u$ has been utilized in crafting the system prompt that the LLM RecSys has used to improve the relevance of recommended items, i.e., to find out whether $u\in\{u_{1},\dots,u_{k}\}$.
4.2 Adversary’s Capabilities
The adversary can access the large language model for RecSys, customized via prompts in a known format, as indicated in previous research in LLM RecSys [25, 30]. The adversary has known the target user’s historical interactions and wanted to know whether they are used in prompts. We consider the most strict and realistic scenario, where the adversary has only black-box access to the target language model $\mathcal{M}$ and its recommended items, but not the tokenizer or the associated output probabilities. We also assume the adversary can access general word embeddings obtained via open-source LLMs (not the target LLM), which can be used in the similarity attack.
5 ATTACK METHODS
5.1 Direct Inquiry Attack
Intuition. The core concept of this attack method hinges on the language model’s ability to remember information from past conversations and deliver context-based responses. When we interact with a language model, it processes the context and produces a response informed by the knowledge it has acquired from previous inputs by the user, particularly from the interaction examples included in the RecSys system prompts. Consequently, an intuitive approach is to directly question the language model about its previous encounters with specific samples.
Figure 4: The direct inquiry attack
Method. The attack methodology is structured as follows (refer to Figure 4 for an illustration):
  • The adversary selects a target user $u$ to determine whether the user showed up in prompts.
  • The adversary crafts a query to the model with the prompt: “Have you seen the user $u$? Only Answer Yes or No”.
  • The adversary sends the query to the model and observes the model’s response. If the model confirms with a “yes”, the user is considered as a member of the dataset; if not, it is considered a non-member.
For LLMs that memorize the prompts, this attack might work. We have witnessed Llama-3 behaves so in experiments.
5.2 Hallucination Attack
Intuition. Hallucination is a common problem with LLMs. In LLM-based RecSys, hallucination may cause the LLM to recommend items out of the domain, i.e., the global item set $I$. When we interact with a language model, it processes the context and recommends items provided by the user conversation and system prompts. We hypothesize that if the model has seen something related to the user before, the probability of hallucinated items might be lower. Specifically, if the number is higher than a threshold, it is likely a non-member.
Figure 5: The hallucination attack
Method. The attack method consists of the following steps (see Figure 5 for an illustration):
  • The adversary selects a target user $u$ to determine the membership.
  • The adversary crafts a query to the model with a prompt, e.g., “Please recommend top-10 movies sorted in descending order of relevance for $u$. Only give movie names without any description”.
  • The adversary counts the number of hallucinated items in the set of recommended items out of the defined item set $I$. If this number is larger than threshold it is considered a non-member, vice versa.
In experiments, we have used IMDB to generate the global item set, $I$, for movies, and MusicBrainz for musics. A threshold of 2 for 10 recommended items seems to work for the experimented models. However, the attack advantages are low.
5.3 Similarity Attack
Intuition. This attack uses the memorization capability of language models to estimate whether the LLM recommended items similar to the user’s historical item interactions known by the adversary. We hypothesize that if the LLM has seen the user’s historical interactions, the recommended items might be similar to historical interactions. This similarity-based attack is borrowed from the attack on traditional RecSys models [42]. However, the similarity calculation is the key. The previous attack made the strong assumption that the adversary also knows the item embedding vectors derived from a large set of known interactions. Considering it is almost impossible to get such embedding vectors in realistic attacks without breaking into the RecSys internal. We redesigned the similarity measuring method for LLM RecSys with general semantic text embedding generated by LLMs. We then use the following method to estimate the similarity between the recommended items and the user’s historical interactions.
Figure 6: The similarity attack
Method. The attack method consists of the following steps (see Figure 6 for an illustration):
  • The adversary selects a target user $u$ to determine its membership status.
  • The adversary crafts a query to the model with a prompt like “Please recommend top-10 movies sorted by a descending order of relevance for $u$. Only give movie names without any description”.
  • The attacker calculates the pairwise similarity between each recommended item and each item in the historical interaction set. The average similarity is used to infer membership status. If the average similarity exceeds a predefined threshold $\tau$, the interaction is classified as a member; otherwise, it is considered a non-member. Formally, let $R=\{r_{1},r_{2},\dots,r_{m}\}$ denote the set of recommended items and $I=\{i_{1},i_{2},\dots,i_{n}\}$ denote the historical interacted items. Let $\text{sim}(r,i)$ denote a similarity function between items $r$ and $i$. The average similarity is computed as:
    $\text{AvgSim}(R_{u},I_{u})=\frac{1}{m\cdot n}\sum_{r\in R_{u}}\sum_{i\in I_{u}}\text{sim}(r,i).$ (1)
    In experiments, we have used the Sentence-Transformer network [34], a widely utilized text encoder, to embed items and the cosine similarity for pairwise similarity calculation. However, we noticed that LLMs will not duplicate the seen historical items or fine-tune the list, even though the user has been included in prompts, which is expected for a reasonable RecSys. Therefore, since the general text embedding does not contain the unique user-item interaction information as the traditional item embedding does [42], we have seen the similarity attacks does not work ideally.
5.4 Poisoning Attack
Intuition. We design the poisoning attack to further exploit the memorization capability of LLMs. We hypothesize that if the model has previously seen the target user’s interactions, it will exhibit a certain degree of “stubbornness”. Specifically, if the adversary shows additional prompts that contain the targeted user’s modified historical interactions, the LLM having the memory of the correct historical interactions is less likely to change its mind. In contrast, if the model has not seen the historical interactions, its recommended items might be more influenced by the provided modified items.
Figure 7: The poisoning attack
Method. It consists of the following steps (see Figure 7 for an illustration):
  • The adversary selects a target user $u$ to determine its membership status.
  • The adversary crafts a query to the model with a prompt like “Please recommend top-10 movies sorted by the descending order of relevance for $u$. Only give movie names without any description.” The system returns the initial recommended items $R_{0}$.
  • The adversary further provides a prompt with the modified historical interactions, e.g., “The user James Wiseman previously interacted with items ($i_{1},i_{2},\dots,i_{n}$), can you recommended another list of 10 movies?” The modified list ($i_{1},i_{2},\dots,i_{n}$) is generated as follows. The adversary randomly selects and replaces items in a user’s original interaction set $I_{u}$ with low-similarity items from the total set $I$. In the movie recommendation case, we used all the movies in the IMDB dataset. Specifically, for a selected item $i^{0}_{k}\in I_{u}$, the attacker substitutes it with an item $i_{k}$ such that:
    $i_{k}=\arg\min_{j\in\text{IMDB}}\text{sim}(i^{0}_{k},j).$ (2)
  • The attacker gets another list of recommended items, $R_{1}$. Then the similarity between $R_{0}$ and $R_{1}$ is calculated with the same average similarity formula Eq. 1. The membership inference decision is based on the similarity gap. If the similarity gap is less than the predefined $\tau$, it implies the LLM’s decision is less likely influenced by the modified list, and the user’s interactions may have shown up in the system’s prompts.
6 EXPERIMENTS
The experiments try to answer the following questions in our attack design. (1) Will the RecSys LLM likely leak the user’s information via direct inquiry? (2) Can the previously reported similarity-checking attack [42, 46] also work in the LLM context? (3) Can the LLM’s memorization feature be explored to infer membership? (4) What are the factors affecting the performance of membership inference?
6.1 Experiment setup
Large Language Models. We evaluate our attacks on three representative language models, including Llama-2 [36], Llama-3 [36], and Vicuna [12] that have been used in studying LLM RecSys [25, 44]. For Llama-2 and Llama-3, we utilize the 7B and 8B versions, respectively. Vicuna is an instruct-finetuned version of Llama-2. We utilize the 13B version of Vicuna. In general, newer models may have improved some aspects, i.e., the reduced hallucinations and better relevance of the recommended items. We use these models for experiments because they have shown their effectiveness for RecSys in previous studies. More recent models, e.g., Llama-4, may show better recommendation performance. However, to isolate the unknown factors, we decide to use these models only for easier comparison and better reproducibility.
Dataset #Users #Items #Interactions
MovieLens-1M 6,040 3,706 1,000,209
Amazon Digital Music 840,372 456,992 1,584,082
Table 1: Statistics of datasets.
Datasets. We assess the impact of our attacks on MovieLens-1M [21] and Amazon Digital Music [26]. The statistic of the dataset is shown in Table 1. For hallucination attack, we use IMDB [1] on the MovieLens-1M as the corpus to decide whether LLM’s recommended item is hallucinated item, which aligned with the previous research [25]. Also, we find the MusicBrainz [2] can be used as corpus on the Amazon Digital Music to decide whether the LLM’s output is hallucinated item. The recommendation-specific prompts are designed using the template provided by Liu et al. [30] known for its verified performance. It is worth noting that our objective is to determine if the target user information is used in the system prompt. However, it is likely that the LLM has seen relevant text samples in the original LLM pre-training data, which we cannot distinguish. However, as demonstrated in this study [39], LLMs exhibit significantly weaker memorization about pre-training data, compared to the data in prompts.
Evaluation Metrics. Since we focus on whether each attack works and their relative performance over different settings, we consider the widely adopted metric in related studies [39] – the attack advantage:
$\text{Adv}=2\times(\text{Acc}-0.5),$ (3)
where Acc is the attack accuracy. It measures the advantage over random guessing: random guessing remains at 0, while a perfect attack gives 1. Other MIA studies [5, 46] have used more fine-grained metrics, such as true positive rate (TPR) at low false positive rate (FPR), to precisely compare different MIA attacks. However, as the first MIA study on LLM RecSys, our focus is to show whether there are effective MIA attacks on the LLM RecSys approach. Thus, attack advantage is sufficient.
Experiment Design. Each dataset is de-duplicated and randomly partitioned into two disjoint subsets: the member set and the non-member set. For each trial, a number of samples, i.e., the “shots” are randomly drawn from the member set to form the system prompt, one of which is selected as the member sample, while a random sample from the non-member set serves as the non-member. Repeating this process 500 times results in a balanced evaluation set consisting of 500 member and 500 non-member samples. We also simulate the RecSys system prompts with a different number of shots to investigate the relationship between the system prompt setting vs attack effectiveness. The 500 independent experiments for each setting, i.e., the specific LLM and the number of shots, ensure the statistical significance of our findings.
6.2 Result Analysis
Our experimental results show that the two attacks: direct inquiry and poisoning attacks work effectively, while the other two: hallucination and similarity attacks are unsatisfactory. We organize them in the same order.
Meanwhile, we found that several factors affect the effectiveness of the attacks, including the number of shots used by the system prompt, the position of the attacked shot in the system prompt, and different LLMs. For each attack, we will explore these important factors.
6.2.1 Direct Inquiry
We find Direct Inquiry works surprisingly well on Llama-3, which likely memorizes and utilizes the prompts better than older models. Older models, Llama-2 and Vicuna, do not work in this attack. Figure 8 shows the comparison on the most effective case: one-shot system prompt. For other settings, the older models give almost zero attack advantage and thus ignored. Both older models tend to blindly give the same answer regardless of whether the sample is a member or nonmember. As this attack fully depends on the basic generative ability and prompt memorization of the LLM, it makes sense that newer LLMs may be more vulnerable to this attack.
Llama-3VicunaLlama-200.50.511Attack AdvantageMovieMusic
Figure 8: Best Direct Inquery performance for each LLM and dataset.
Figure 9 shows more details about this attack on both datasets with Llama-3. The attack performance is also affected by the number of shots used by the system prompt, and the shot position that contains the targeted user. Interestingly, the more shots used by the system prompt, the less likely this attack will succeed, which implies that the system designer can simply increase the shots to make the RecSys more resilient to this type of attack. The targeted shot position also plays a role in attack results. “Movie-first” in the figure means the victim happens to be at the first position of the system prompt. The victim at the earlier shot positions seems more vulnerable than later ones.
1-shot5-shot10-shot00.50.511Attack AdvantageMovie-firstMovie-lastMusic-firstMusic-last
Figure 9: Attack advantage of direct inquiry against Llama-3 with different shot counts and attacked shot positions (e.g., Movie-first for the first shot position on the movie dataset).
6.2.2 Poisoning Attack
Poisoning attack shows different patterns from the direct inquiry attack. This attack has a unique additional factor: the number of poisoned items, which are crafted by the attacker in the additional attacker-provided prompt. We find that both the number of shots in system prompts and the number of poisoned items in attacker’s prompts affect the attack performance. Interestingly, different datasets may show different affecting patterns, possibly due to the dataset-related topic and its relevant content in the original pre-training dataset. Similar to the first experiment, we experimented with 1-shot, 5-shot, and 10-shot system prompts. In addition, we investigated 2-poison, 5-poison, and 8-poison items in the attacker’s prompts. With a similarity gap threshold $\tau$ around 0.3 to 0.45, we can achieve satisfactory MIA performance. Figure 10 summarizes the best performing MIA result under all settings for each LLM and each dataset. Different from the previous attack, an old-version LLM, Llama-2, shows the best performance for both datasets.
Llama-2Llama-3Vicuna00.50.511Attack AdvantageMovieMusic
Figure 10: Best attack advantage under poisoning for different models and domains
We look into the Llama-2 results in more detail. Since the poisoning process randomly takes a number of items to poison, we do not distinguish the effect of poisoning the first or the last shot. Figure 11 on the movie dataset shows that more poisoned items may slightly increase the chance of MIA success, while more shots may reduce the MIA performance. Figure 12 on the music dataset shows that the number of shots has more significant effect on MIA: the more shots, the worse MIA performs, while the pattern of the number of poisoned items is unclear.
1-shot5-shot10-shot0.60.60.80.811Attack Advantage2-poisoned5-poisoned8-poisoned
Figure 11: Poisoning Attack advantage for Llama-2 on movie dataset with k-shots in the system prompt and m number of poisoned items provided by the attacker. The advantage increases with more poisoned items.
1-shot5-shot10-shot00.20.20.40.40.60.60.80.811Attack Advantage2-poisoned5-poisoned8-poisoned
Figure 12: Poisoning Attack advantage for Llama-2 on music dataset with k-shots in the system prompt and m number of poisoned items. The advantage decreases with more shots
6.2.3 Hallucination Attack
This attack performs much worse than the previous two. The best performing is the Vicuna model on the music data with 5-shot system prompt, around 0.41 advantage. Most other settings give advantages around 0.1 or lower. We take the threshold = 2, i.e., the number of hallucinated items, to classify members and non-members, which gives the best performance. Figure 13 summarizes the best results for each dataset and each LLM on 1-shot, 5-shot, and 10-shot settings.
VicunaLlama-2Llama-300.20.20.40.4Attack AdvantageMovieMusic
Figure 13: Vicuna is the most vulnerable in the hallucination attack. However, this attack has significantly worse performance than the direct inquiry and the poisoning attacks.
Figure 14 looks into the details for the best performing Vinuna model. We also examine whether the attacked shot position (first and last) affects the result. While the attack performance was miserable for the movie dataset, it does show a trend of slight increasing performance with the increase of shots. However, there is no clear pattern on attacking the first or the last shot position, although the best performing attack happens at the last shot on the 5-shot setting for the music dataset.
1-shot5-shot10-shot00.20.20.40.4Attack AdvantageMovie-firstMovie-lastMusic-firstMusic-last
Figure 14: Hallucination attack with Vicuna. There is an increasing trend with the increase of shots for the movie dataset, while there is no clear pattern about the attacked shot positions.
6.2.4 Similarity Attack
Similarity attack is so far the worst performing one. Figure 15 shows the best result for each LLM and dataset. Llama-2 performs the best on the movie dataset, but only achieves around 0.23 advantage. The second best is Vicuna on movie, around 0.15. The other items are slightly above 0.
Llama-2VicunaLlama-300.10.10.20.20.30.3Attack AdvantageMovieMusic
Figure 15: Llama-2 is the most vulnerable in the similarity attack. However, this attack overall performs the worst among other attacks.
The similarity attack on traditional RecSys [42] utilizes the item embeddings derived from the interaction matrix to achieve good attacking performance. While such a known RecSys-specific embedding scheme is a very strong assumption, our poor similarity attack performance indicates that the general text embedding scheme may not have the information that the interaction-based embedding has.
6.3 Discussion
In our initial study of the novel MIAs on LLM-based RecSys, we have identified that at least two effective attacks: direct inquiry and poisoning. They perform surprisingly well on the well-known public RecSys datasets. However, the use of different versions of LLMs leads to diverse results, which is partially tied to the match between the nature of the attack and the features of different versions of LLMs. For example, the direct inquiry attack utilizes the prompt memorization feature and the generalization power of the LLMs, the later models may give better performance. However, this type of attacks can be easily mitigated via model alignment, i.e., fine-tuning the model not to answer such kind of queries. The poisoning attack somehow utilizes the weakness, i.e., the memorization ability of older LLMs and thus these LLMs are more vulnerable to this attack. Recent LLMs have likely incorporated various techniques [20, 35] to address the memorization issue and thus less influenced by such attacks.
The hallucination attack performs unsatisfactorily. Our results were almost against our initial design hypothesis: the seen information may lead to less hallucination. This indeed happened for a few cases, e.g., older models, like Llama-2 that likely has stronger memorization and hallucination, but the attack advantage is marginal.
The similarity attack performs badly. It is partly due to the significant difference between the general text embedding and the interaction-matrix based embedding. However, it also leaves space for improvement. There might be a better similarity attack to achieve performance comparable to our top ranked attacks.
7 Conclusion
With LLMs potentially used for developing next-generation RecSys, the new privacy risks should be thoroughly understood. We designed four novel MIA attacks, specifically targeting LLM-based RecSys, and have shown that at least two of the attacks: direct inquiry and poisoning work effectively. Thus, we confirm that the membership inference threat on LLM-based RecSys is realistic. Practitioners and researchers of LLM RecSys should carefully design schemes to address such privacy threats. We will also extend this study to more MIA attacks and design corresponding mitigation methods.